Calculate the efficiency of an ice making machine that takes in water at 10°C and produces ice at a temperature of -7°C. Water is taken in at a rate of 1 kg every 4 minutes. The input power of the machine is 400 W. 

 

The specific heat capacity of water is 4187 Jkg−1K−1 and its specific latent heat of fusion is 334 Jkg−1.


1
Expert's answer
2021-05-05T07:46:30-0400

Mass of flow rate of water = Rate of ice making(m)="\frac{1 kg}{4\times 60 sec}" =4.166"\times 10^{-3}" kg/sec

Amount of heat absorbed (Q)= Sensible Heat +latent heat=m(Cp"\Delta T+h_{fg}" )

=4.166"\times 10^{-3}" (4187x"17^o+334" )=296.48 watt

Input Power(W)=400 watt

Efficiency of cooling system ="\frac{}{}" "\frac {Desired\: Cooling \: Effect}{Input \: Power}\times 100=\frac{296.48}{400}\times 100" =73.21 %